# Collaborative Coded Caching for Partially Connected Networks

### v2

@@ Abstract
Coded caching leverages the differences in user cache memories to achieve gains that scale with the total cache size, alleviating network congestion due to high-quality content requests. Additionally, distributing transmitters over a wide area can mitigate the adverse effects of path loss. In this work, we consider a partially connected network where the channel between distributed transmitters (helpers) and users is modeled as a distributed MIMO Gaussian broadcast channel. We propose a novel delivery scheme consisting of two phases: partitioning and transmission. In the partitioning phase, users with identical cache profiles are partitioned into the minimum number of sets, such that users within each set can successfully decode their desired message from a joint transmission enabled by MIMO precoding. To optimally partition the users, we employ the branch and bound method. In the transmission phase, each partition is treated as a single entity, and codewords are multicast to partitions with distinct cache profiles. The proposed delivery scheme is applicable to any partially connected network, and while the partitioning is optimal, the overall delivery scheme, including transmission, is heuristic. Interestingly, simulation results show that its performance closely approximates that of the fully connected optimal solution.
@@ I Introduction
To address these challenges, various technologies have been proposed to mitigate the impact of the limitations above. For example, to counteract the effects of path loss, one promising solution involves the deployment of multiple transmitters, such as WiFi routers, spread across a wide area [1]. This ensures that users remain within a reasonable distance from at least one transmitter, improving signal strength and reducing the likelihood of experiencing low data rates.
Meanwhile, the demand for content-related applications has been steadily increasing, with corresponding quality requirements rising. One promising technology that has the potential to alleviate these challenges is coded caching [2]. With coded caching, users can pre-store some of the file contents in off-peak times to reduce network congestion and achieve gains scaled with the cumulative cache size of the network. Many works have built on the original coded caching work of [2], e.g., for multi-server [3], Device-to-Device (D2D) [4], transmitter-side cache [5], shared-cache [6], multi-access [7], combinatorial [8], dynamic [9] networks, and  [10, 11, 12] combined multi-antenna/transmitter strategies with coded caching.
Figure 1: A partially connected network with $R=2$ transmitters/helpers, $K=7$ users and $L=3$ cache profiles.
@@
This paper builds upon the shared-cache model of [6] to reduce subpacketization, where some users share the same cache placement. Inspired by [11, 12], we leverage the coded caching gain to serve users with different cache profiles, and the spatial multiplexing gain of the transmitters to serve users with identical cache profiles simultaneously. Unlike [11, 12], where each user is within the coverage area of all transmitters, we face the challenge that each user is only associated with a limited subset of transmitters. The fully connected network model is relevant in a single-cell scenario with a MIMO base station, where all users are within reach of all antennas. In contrast, in an extended network where helpers are spatially distributed, it is unrealistic to assume that each user can receive a sufficiently strong signal from all helpers. This motivates the use of the partially connected geometric model, as proposed in [13, 14, 15], while also considering full cooperation among all helpers at the physical layer through MIMO precoding. In a fully connected network, any subset of users with the same cache profile, equal in number to the transmitters or antennas, can be served together as in [11, 12]. However, in a partially connected network, the subset of users with the same cache profile that can be served together depends on the specific network topology.
To address this, we propose a novel delivery scheme that combines coded caching and spatial multiplexing to optimize data delivery in partially connected networks. The proposed scheme consists of two parts: partitioning and transmission. In the partitioning phase, we exploit the spatial multiplexing gain of the transmitters to partition users with the same cache profile into the minimum number of sets, such that users within each set can successfully decode their desired message from a joint transmission. To the best of our knowledge, this partitioning problem has not been previously studied. To solve it optimally, we employ the least cost branch and bound method, which is commonly used to solve combinatorial optimization problems [16, 17, 18]. In the transmission phase, we leverage the coded caching gain by treating each partition of users as a single entity. Codewords are then multicast to different partitions, each with a distinct cache profile, following the approach in [11, 12].
@@ II System Model
In our model, a server is connected to $R$ single antenna helpers (transmitters) via error-free fronthaul links and serves the requests of $K$ cache-enabled users. We assume that users are interested in receiving files from a library of $N$ files and can store a fraction of this library in their cache memory. Each library file is $F$ bits, and each user can cache $MF$ bits corresponding to a fraction $\gamma=M/N$ of all the files. We assume that the fronthaul links have sufficiently high capacity so they do not form the communication bottleneck and each helper has an effective transmission radius, $r$, such that any links between a helper and users outside of this radius are equal to zero. In contrast, the links for users within the radius are non-zero.
@@
where ${\mathbf{x}}\in\mathbb{C}^{R\times 1}$ denotes the vector containing all the signals transmitted by the $R$ helpers satisfying an average power constraint $\mathbb{E}(||{\mathbf{x}}||^{2})\leq P$, ${\mathbf{h}}_{k}\in\mathbb{C}^{R\times 1}$ denotes the channel between each of the $R$ helpers and user $u_{k}$ following some i.i.d continuous distribution (e.g. Gaussian) where ${\mathbf{h}}_{k}[i]$=0 if user $u_{k}$ is outside of the transmission radius of helper $h_{i}$, and $w_{k}$ represents the unit power AWGN noise at user $u_{k}$. We assume that perfect channel state information exists throughout the (active) nodes as in [3, 5, 11, 12], the fading process is statistically symmetric across users where each non-zero link between a helper and a user has a capacity of the form $\log(SNR)+o\log(SNR)$, and the system operates in a very high power regime such that the metric of interest is the sum-Degrees of Freedom (DoF), i.e., the number of users that can be served simultaneously with a rate that grows as $\log(P)$ for asymptotically large $P$.
A time slot corresponds to transmitting a single file from the library to a single user, and we want to minimize the worst-case delivery time T, where T denotes the number of time slots required to serve all the users when each user requests a distinct file from the library, given a cache placement. In the seminal coded caching paper [2], each file in the library is split into $\binom{K}{K\gamma}$ sub-files such that the caches of the users are filled with collections of these subfiles where the cache of each user is unique. The users having different caches is significantly important since codewords can be multicast to multiple users using the differences in caches where the exact number of users to whom codewords can be multicast turns out to be $K\gamma+1$. However, since $\binom{K}{K\gamma}$ grows exponentially with $K$ in the placement scheme of [2], it requires a more practical approach for finite-sized files.
@@ III Description of the Scheme
In this paper, we employ the cache placement scheme of [6], where users are assigned to $L<K$ distinct cache profiles, and the cache content of each user is given by its assigned profile. The main advantage of this cache placement scheme is its reduced subpacketization requirement [23, 24, 25, 26, 27, 28], and its practicality, as displayed in  [12, 13, 14, 15].
@@
The network of interest is partially connected. A user is within the transmission radius of one or multiple transmitters. With coded caching, we can serve multiple users with different cache profiles at the same time, and by using the spatial multiplexing gain of the transmitters, we can also serve the users with the same cache profile together. This idea is inspired by the schemes of [11, 12], where their delivery schemes are stated in the following:
Let ${\mathbf{c}}_{\ell}$ denote the vector of user indices assigned to cache profile $\ell$ with size $|c_{\ell}|=C_{\ell}$, and $d_{{\mathbf{c}}_{\ell}[j]}$ denote the index of the file requested by user ${\mathbf{c}}_{\ell}[j]$, for $j\in[C_{\ell}]$. For a fully connected network where there is a nonzero link between each helper (transmitter) and each user, and $C_{\ell}=R$ for each $\ell\in[L]$, in [11] the following vector
@@
is transmitted for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$ to satisfy all the user requests where ${\mathbf{H}}_{({\mathbf{c}}_{\ell})}^{-1}$ denotes the inverse of the channel matrix between the $R$ transmitters and the $C_{\ell}=R$ users with cache profile $\ell$, and ${\mathbf{w}}_{({\mathbf{c}}_{\ell})}$ denotes the column vector where its $j$th element is $\𝒯W^{d_{{\mathbf{c}}_{\ell}[j]}}_{{\mathcal{T}}\backslash\{\ell\}}$, for $j\in[C_{\ell}]$. So, in each transmission $R(\gamma L+1)$ users are served simultaneously. After (2) is transmitted, user $u_{k^{\prime}}$ with cache profile $\ell^{\prime}$ where $\ell^{\prime}\in{{\mathcal{T}}}$ can remove $\sum_{\ell\in{\mathcal{T}}\backslash\{\ell^{\prime}\}}{\mathbf{h}}_{k^{\prime}}^{T}{\mathbf{H}}_{({\mathbf{c}}_{\ell})}^{-1}{\mathbf{w}}_{({\mathbf{c}}_{\ell})}$ from its received message since $\𝒯\ell^{\prime}\in{{\mathcal{T}}\backslash\{\ell\}}$ and $\𝒯W^{d_{{\mathbf{c}}_{\ell}[j]}}_{{\mathcal{T}}\backslash\{\ell\}}$ is in the cache of user $u_{k^{\prime}}$, for every $\ell\in{\mathcal{T}}\backslash\{\ell^{\prime}\}$ and every $j\in[C_{\ell}]$. User $u_{k}{{}^{\prime}}$ then can decode its desired subfile $W^{d_{{\mathbf{c}}_{\ell^{\prime}}[j^{\prime}]}}_{{\mathcal{T}}\backslash\{\ell^{\prime}\}}$, where $j^{\prime}$ corresponds to the index of $u_{k^{\prime}}$ in the vector of users with cache profile $\ell^{\prime}$, since ${\mathbf{h}}_{k^{\prime}}^{T}{\mathbf{H}}_{({\mathbf{c}}_{\ell^{\prime}})}^{-1}{\mathbf{w}}_{({\mathbf{c}}_{\ell^{\prime}})}=W^{d_{{\mathbf{c}}_{\ell^{\prime}}[j^{\prime}]}}_{{\mathcal{T}}\backslash\{\ell^{\prime}\}}$. In this way, each user with cache profile $\ell\in{{\mathcal{T}}}$ can decode a subfile from each transmission such that after (2) is transmitted for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$, each user can decode their requested file.
@@
. For a fully connected network where there is a nonzero link between each helper (transmitter) and each user with $C_{\ell}\geq R$ for each $\ell\in[L]$[12] further divides each subfile $W^{n}_{{\mathcal{S}}}$ into $R$ equal-sized smaller subfiles $W^{n}_{{\mathcal{S}},i}$, for $i\in[R]$. User requests are satisfied in $\max_{\ell}{C_{\ell}}:=C$ rounds. Let ${\mathbf{c}}_{\ell}^{R}:=[{\mathbf{c}}_{\ell},{\mathbf{c}}_{\ell},...,{\mathbf{c}}_{\ell}]$ denote the $R$-fold concatenated vector of ${\mathbf{c}}_{\ell}$ for $\ell\in[L]$. In each round $o$, $o\in[C]$, users ${\mathbf{c}}_{\ell}^{R}[(o-1)*R+i]$ for each $i\in[R]$ and each $\ell\in[L]$ are served, if the element of the vector is nonempty. Notice that for given $o$ and $\ell$, ${\mathbf{c}}_{\ell}^{R}[(o-1)*R+i]$ is either empty or nonempty for all $i\in[R]$. For given ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=t+1$, let ${\mathcal{T}}^{o}$ denote the set ${\mathcal{T}}$ excluding the cache profiles $\ell$ where ${\mathbf{c}}_{\ell}^{R}[(o-1)*R+i]$ is empty for all $i\in[R]$ in round $o$. Then in round $o$, for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$ where ${\mathcal{T}}^{o}$ is nonempty, the following vector is transmitted
@@
where ${\mathbf{H}}_{({\mathbf{c}}_{\ell}^{R}(o))}^{-1}$ denotes the inverse of the channel matrix between the $R$ transmitters and the $R$ users with cache profile $\ell$ in round $o$, and ${\mathbf{w}}_{({\mathbf{c}}_{\ell}^{R}(o))}$ denotes the column vector where its $i$th element is $W^{d_{{\mathbf{c}}_{\ell}^{R}[(o-1)*R+i]}}_{{\mathcal{T}}\backslash\{\ell\},(o,i)}$ for $i\in[R]$, $(o,i)$ corresponds to the next subindex of the subfile requested by user ${\mathbf{c}}_{\ell}^{R}[(o-1)*R+i]$. Decoding follows similarly as the decoding for the transmission in (2).
[11, 12] consider a fully connected network and assume that the channels between each set of $R$ users and the $R$ transmitters are invertible. In this way, they can serve any $R$ users with the same cache profile using MIMO precoding, and their delivery schemes are based on this simple idea. However, for a partially connected network with possibly zero links between some transmitters and some users, as considered in this paper, this can not be assumed so that the schemes in [11, 12] are not applicable. The problem is assigning the users to transmitters; since not every user can be assigned to each helper, if a user has multiple helpers that can be assigned to, then which helper the user is assigned to may affect the performance of a delivery scheme. In this paper, we propose a delivery scheme with two parts: partitioning and transmission. In partitioning, users corresponding to each cache profile are partitioned into a minimum number of sets where each user in a set can successfully decode their desired message from the joint transmission dedicated to the set. After users are partitioned accordingly, in transmission, each partition of users is treated as a single user, and codewords are multicasted to different partitions with different cache profiles as in [11, 12].

### v3

@@ Abstract
Coded caching leverages the differences in user cache memories to achieve gains that scale with the total cache size, alleviating network congestion due to high-quality content requests. Additionally, distributing transmitters over a wide area can mitigate the adverse effects of path loss. In this work, we consider a partially connected network where the channel between distributed transmitters (helpers) and users is modeled as a distributed multiple-input-multiple-output (MIMO) Gaussian broadcast channel. We propose a novel delivery scheme consisting of two phases: partitioning and transmission. In the partitioning phase, users with identical cache profiles are partitioned into the minimum number of sets, such that users within each set can successfully decode their desired message from a joint transmission enabled by MIMO precoding. To optimally partition the users, we employ the branch and bound method. In the transmission phase, each partition is treated as a single entity, and codewords are multicast to partitions with distinct cache profiles. The proposed delivery scheme is applicable to any partially connected network, and while the partitioning is optimal, the overall delivery scheme, including transmission, is heuristic. Interestingly, simulation results show that its performance closely approximates that of the fully connected optimal solution.
@@ I Introduction
To address these challenges, various technologies have been proposed to mitigate the impact of the limitations above. For example, to counteract the effects of path loss, one promising solution involves deploying multiple transmitters, such as WiFi routers, spread across a wide area [1]. This ensures that users remain within a reasonable distance from at least one transmitter, improving signal strength and reducing the likelihood of experiencing low data rates.
Meanwhile, the demand for content-related applications has been steadily increasing, along with rising quality requirements. One promising technology that has the potential to alleviate these challenges is coded caching [2]. With coded caching, users can pre-store some of the file contents during off-peak times to reduce network congestion and achieve gains that scale with the cumulative cache size of the network. Many works have built on the original coded caching work of [2], e.g., for multi-server [3], Device-to-Device (D2D) [4], transmitter-side cache [5], shared-cache [6], multi-access [7], combinatorial [8], dynamic [9] networks, and  [10, 11, 12] combined multi-antenna/transmitter strategies with coded caching.
Figure 1: A partially connected network with $E=2$ transmitters/helpers, $K=7$ users and $L=3$ cache profiles.
@@
This paper builds upon the shared-cache model of [6] to reduce subpacketization, where some users share the same cache placement. Inspired by [11, 12], we leverage the coded caching gain to serve users with different cache profiles, and the spatial multiplexing gain of the transmitters to serve users with identical cache profiles simultaneously. Unlike [11, 12], where each user is within the coverage area of all transmitters, we face the challenge that each user is only associated with a limited subset of transmitters. The fully connected network model is relevant in a single-cell scenario with a multi-antenna base station, where all users are within reach of all antennas. In contrast, in an extended network where helpers are spatially distributed, it is unrealistic to assume that each user can receive a sufficiently strong signal from all helpers. This motivates using the partially connected geometric model, as proposed in [13, 14, 15], while also considering full cooperation among all helpers at the physical layer through MIMO precoding. In a fully connected network, any subset of users with the same cache profile, equal in number to the transmitters or antennas, can be served together, as in [11, 12]. However, in a partially connected network, the subset of users with the same cache profile that can be served together depends on the specific network topology.
To address this, we propose a novel delivery scheme that combines coded caching and spatial multiplexing to optimize data delivery in partially connected networks. The proposed scheme consists of two parts: partitioning and transmission. In the partitioning phase, we exploit the spatial multiplexing gain of the transmitters to partition users with the same cache profile into the minimum number of sets, such that users within each set can successfully decode their desired message from a joint transmission. To the best of our knowledge, this partitioning problem has not been previously studied. To solve it optimally, we employ the least cost branch and bound method, which is commonly used to solve combinatorial optimization problems [16, 17, 18]. In the transmission phase, we leverage the coded caching gain by treating each user partition as a single entity. Codewords are then multicast to different partitions, each with a distinct cache profile, following the approach in [11, 12].
@@ II System Model
In our model, a server is connected to $E$ single antenna helpers (transmitters) via error-free fronthaul links and serves the requests of $K$ cache-enabled users. We assume that users are interested in receiving files from a library of $N$ files and can store a fraction of this library in their cache memory. Each library file is $F$ bits, and each user can cache $MF$ bits corresponding to a fraction $\gamma=M/N$ of all the files. We assume that the fronthaul links have sufficiently high capacity, so they do not form the communication bottleneck and each helper has an effective transmission radius, $r$, such that any links between a helper and users outside of this radius are equal to zero. In contrast, the links for users within the radius are nonzero.
@@
where ${\mathbf{x}}\in\mathbb{C}^{E\times 1}$ denotes the vector containing all the signals transmitted by the $E$ helpers satisfying an average power constraint $\mathbb{E}(||{\mathbf{x}}||^{2})\leq P$, ${\mathbf{h}}_{k}\in\mathbb{C}^{E\times 1}$ denotes the channel vector where ${\mathbf{h}}_{k}[i]$ is the channel coefficient between helper $e_{i}$ and user $u_{k}$ following some i.i.d continuous distribution (e.g. Gaussian) and is 00 if user $u_{k}$ is outside of the transmission radius of helper $e_{i}$, and $w_{k}$ represents the unit power AWGN noise at user $u_{k}$. We assume that perfect channel state information exists throughout the (active) nodes as in [3, 5, 11, 12], the fading process is statistically symmetric across users where each nonzero link between a helper and a user has a capacity of the form $\log(SNR)+o(\log(SNR))$, and the system operates in a very high power regime such that the metric of interest is the sum-Degrees of Freedom (DoF), i.e., the number of users that can be served simultaneously with a rate that grows as $\log(P)$ for asymptotically large $P$.
A time slot corresponds to transmitting a single file from the library to a single user, and we want to minimize the worst-case delivery time T, where T denotes the number of time slots required to serve all the users when each user requests a distinct file from the library, given a cache placement. In this way, the sum-DoF becomes $d_{\sum}=\frac{K(1-\gamma)}{T}$, when each user caches $\gamma$ portion of each file. In the seminal coded caching paper [2], each file in the library is split into $\binom{K}{K\gamma}$ sub-files such that the caches of the users are filled with collections of these subfiles where the cache of each user is unique. The users having different caches is significantly important since codewords can be multicast to multiple users using the differences in caches where the exact number of users to whom codewords can be multicast turns out to be $K\gamma+1$. However, since $\binom{K}{K\gamma}$ grows exponentially with $K$ in the placement scheme of [2], it requires a more practical approach for finite-sized files.
@@ III Description of the Scheme
In this paper, we employ the cache placement scheme of [6], where users are assigned to $L<K$ distinct cache profiles, and the cache content of each user is determined by its assigned profile. The main advantage of this cache placement scheme is its reduced subpacketization requirement [23, 24, 25, 26, 27, 28], and its practicality, as displayed in  [12, 13, 14, 15].
@@
The network of interest is partially connected. A user is within the transmission radius of one or multiple transmitters. With coded caching, we can serve multiple users with different cache profiles simultaneously. By utilizing the spatial multiplexing gain of the transmitters, we can also serve users with the same cache profile together. This idea is inspired by the schemes of [11, 12]. where their delivery schemes are stated in the following:
Let ${\mathbf{c}}_{\ell}$ denote the vector of user indices assigned to cache profile $\ell$ with size $|{\mathbf{c}}_{\ell}|=C_{\ell}$, and $d_{{\mathbf{c}}_{\ell}[j]}$ denote the index of the file requested by user ${\mathbf{c}}_{\ell}[j]$, for $j\in[C_{\ell}]$. For a fully connected network where there is a nonzero link between each helper (transmitter) and each user, and $C_{\ell}=E$ for each $\ell\in[L]$, in [11] the following vector
@@
is transmitted for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$ to satisfy all the user requests where ${\mathbf{H}}_{({\mathbf{c}}_{\ell})}^{-1}$ denotes the inverse of the channel matrix between the $E$ transmitters and the $C_{\ell}=E$ users with cache profile $\ell$, and ${\mathbf{w}}_{({\mathbf{c}}_{\ell})}$ denotes the column vector where its $j$th element is $\𝒯W^{d_{{\mathbf{c}}_{\ell}[j]}}_{{\mathcal{T}}\backslash\{\ell\}}$, for $j\in[C_{\ell}]$. So, in each transmission, $E(\gamma L+1)$ users are served simultaneously. After (2) is transmitted, user $u_{k^{\prime}}$ with cache profile $\ell^{\prime}$ where $\ell^{\prime}\in{{\mathcal{T}}}$ can remove $\sum_{\ell\in{\mathcal{T}}\backslash\{\ell^{\prime}\}}{\mathbf{h}}_{k^{\prime}}^{T}{\mathbf{H}}_{({\mathbf{c}}_{\ell})}^{-1}{\mathbf{w}}_{({\mathbf{c}}_{\ell})}$ from its received message since $\𝒯\ell^{\prime}\in{{\mathcal{T}}\backslash\{\ell\}}$ and $\𝒯W^{d_{{\mathbf{c}}_{\ell}[j]}}_{{\mathcal{T}}\backslash\{\ell\}}$ is in the cache of user $u_{k^{\prime}}$, for every $\ell\in{\mathcal{T}}\backslash\{\ell^{\prime}\}$ and every $j\in[C_{\ell}]$. User $u_{k}{{}^{\prime}}$ then can decode its desired subfile $W^{d_{{\mathbf{c}}_{\ell^{\prime}}[j^{\prime}]}}_{{\mathcal{T}}\backslash\{\ell^{\prime}\}}$, where $j^{\prime}$ corresponds to the index of $u_{k^{\prime}}$ in the vector of users with cache profile $\ell^{\prime}$, since ${\mathbf{h}}_{k^{\prime}}^{T}{\mathbf{H}}_{({\mathbf{c}}_{\ell^{\prime}})}^{-1}{\mathbf{w}}_{({\mathbf{c}}_{\ell^{\prime}})}=W^{d_{{\mathbf{c}}_{\ell^{\prime}}[j^{\prime}]}}_{{\mathcal{T}}\backslash\{\ell^{\prime}\}}$. In this way, each user with cache profile $\ell\in{{\mathcal{T}}}$ can decode a subfile from each transmission such that after (2) is transmitted for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$, each user can decode their requested file.
@@
For a fully connected network where there is a nonzero link between each helper (transmitter) and each user with $C_{\ell}\geq E$ for each $\ell\in[L]$[12] further divides each subfile $W^{n}_{{\mathcal{S}}}$ into $E$ equal-sized smaller subfiles $W^{n}_{{\mathcal{S}},i}$, for $i\in[E]$. User requests are satisfied in $\max_{\ell}{C_{\ell}}:=C$ rounds. Let ${\mathbf{c}}_{\ell}^{E}:=[{\mathbf{c}}_{\ell},{\mathbf{c}}_{\ell},...,{\mathbf{c}}_{\ell}]$ denote the $E$-fold concatenated vector of ${\mathbf{c}}_{\ell}$ for $\ell\in[L]$. In each round $o$, $o\in[C]$, users ${\mathbf{c}}_{\ell}^{E}[(o-1)*E+i]$ for each $i\in[E]$ and each $\ell\in[L]$ are served, if the element of the vector is nonempty. Notice that for given $o$ and $\ell$, ${\mathbf{c}}_{\ell}^{E}[(o-1)*E+i]$ is either empty or nonempty for all $i\in[E]$. For given ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=t+1$, let ${\mathcal{T}}^{o}$ denote the set ${\mathcal{T}}$ excluding the cache profiles $\ell$ where ${\mathbf{c}}_{\ell}^{E}[(o-1)*E+i]$ is empty for all $i\in[R]$ in round $o$. Then in round $o$, for each subset ${\mathcal{T}}\subseteq[L]$ of size $|{\mathcal{T}}|=\gamma L+1$ where ${\mathcal{T}}^{o}$ is nonempty, the following vector is transmitted
@@
where ${\mathbf{H}}_{({\mathbf{c}}_{\ell}^{E}(o))}^{-1}$ denotes the inverse of the channel matrix between the $E$ transmitters and the $E$ users with cache profile $\ell$ in round $o$, and ${\mathbf{w}}_{({\mathbf{c}}_{\ell}^{E}(o))}$ denotes the column vector where its $i$th element is $W^{d_{{\mathbf{c}}_{\ell}^{E}[(o-1)*E+i]}}_{{\mathcal{T}}\backslash\{\ell\},(o,i)}$ for $i\in[E]$, $(o,i)$ corresponds to the next subindex of the subfile requested by user ${\mathbf{c}}_{\ell}^{E}[(o-1)*E+i]$. Decoding follows similarly to the decoding for the transmission in (2).
[11, 12] consider a fully connected network and assume that the channels between each set of $E$ users and the $E$ transmitters are invertible. In this way, they can serve any $E$ users with the same cache profile using MIMO precoding, and their delivery schemes are based on this simple idea. However, for a partially connected network with possibly zero links between some transmitters and some users, as considered in this paper, this can not be assumed so that the schemes in [11, 12] are not applicable. The problem is assigning users to transmitters. Since not every user can be assigned to each helper, if a user has multiple helpers that can be assigned to, then which helper the user is assigned to may affect the performance of the delivery scheme. In this paper, we propose a delivery scheme with two parts: partitioning and transmission. In partitioning, users corresponding to each cache profile are partitioned into the minimum number of sets where each user in a set can successfully decode its desired message from the joint transmission dedicated to the set. After users are partitioned accordingly, in transmission, each partition of users is treated as a single user, and codewords are multicast to different partitions with different cache profiles, as in [11, 12].